# Physics and Deep Learning in Computational Wave Imaging

### v1

@@ V Computational Wave Imaging Applications
CWI is widely used in various scientific and engineering fields. This section provides examples of employing these methods to solve real-world problems. Appendix E includes source codes, packages, and training datasets, offering practical examples and a foundation for readers to address their own CWI challenges.
@@ V-A1 Leakage Detection for Carbon Sequestration
Geologic carbon sequestration (GCS), crucial in climate change mitigation, involves capturing and storing CO2. Seismic imaging is key for early detection of CO2 leakage, essential for project success. Effective leakage detection hinges on identifying leak presence, location, and magnitude. While traditional seismic methods are commonly used, they often lack early detection precision. Deep learning has brought significant improvements. Zhou et al. [130] develop a method for direct CO2 leakage quantification from seismic data. Rojas et al. [47] introduce a semi-supervised approach targeting small leaks and imbalanced data scenarios. Yang et al. [64] employ a generative model to enhance rare event detectability. Moving beyond indirect inference from seismic data, Zhong et al. [131] use a generative adversarial network for mapping dynamic saturation changes. Leong et al. [132] create a network for direct CO2 saturation mapping from seismic data. Um et al. [133, 110] combine seismic, electromagnetic, and gravity data for a comprehensive saturation model with uncertainty estimation. Feng et al. [49] link seismic and electromagnetic data for improved CO2 saturation modeling.
@@
The Kimberlina dataset, generated by the National Risk Assessment Partnership project of the U.S. Department of Energy, is based on the geologic structure of a commercial-scale GCS reservoir at the Kimberlina site in the southern San Joaquin Basin, CA, USA. The simulation includes four modules: a CO2 storage reservoir model (Fig. 5(a)), a wellbore leakage model (Fig. 5(b)), multi-phase flow and reactive transport models of CO2/brine migration in aquifers (Fig. 5(c)), and a geophysical model. The P-wave velocity models used in this work, part of the geophysical model, are based on geologic-layer properties from the GCS site (Fig. 5(b)) [128, 129]. Fig. 5(d) illustrates a set of simulations with 20 velocity maps over 200 years. Fig. 5(e) visualizes the inverted velocity maps using physics-based seismic FWI [39], a fully-supervised method [41], and semi-supervised learning approaches [47] to characterize CO2 leakage plumes at the Kimberlina site [134]. Physics-based FWI lacks sufficient resolution to capture early-stage leakage plumes. The fully-supervised method distorts the plume shape to some extent, while the semi-supervised approach yields the most accurate results.
@@
Figure 6: Illustration of seismic full-waveform inversion with uncertainty quantification applying to noisy data. Ground Truth (Row 1); Results obtained using InversionNet [41] on clean data (Row 2); Results obtained using InversionNet [41] on noisy data with SNR=10 (Row 3); Results obtained using InvNet_UQ [109] on noisy data with SNR=10 (Row 4); Uncertainty map obtained using Liu et al. ’s work [109] (Row 5).
@@ V-A2 Seismic Full-Waveform Inversion with Uncertainty Quantification
Robustness to noise is essential for CWI and broader inversion methods. It is important to distinguish “adversarial noise”—deliberate perturbations meant to deceive ML models—from “measurement noise,” which naturally occurs during data collection. Since the latter is far more relevant to CWI, we consider adversarial scenarios to be beyond the scope of this review. In our second seismic FWI example, we specifically address measurement noise. To improve ML models’ resistance to this, researchers have developed strategies such as UQ [109], advanced physics-guided regularization[11], and training with noisy data. We focus on the UQ method InvNet_UQ by Liu et al. [109].
The effectiveness of InvNet_UQ is demonstrated in Fig. 6, where it is applied to data contaminated with noise at a Signal-to-Noise Ratio (SNR) of 10. This is compared to results from InversionNet [41], revealing a significant performance decline for InversionNet under noisy conditions. In contrast, InvNet_UQ maintains higher accuracy in its predictions, showcasing its superior robustness to noise. Furthermore, the uncertainty maps shown in the bottom row of Fig. 6, calculated using Eq. (15), not only underscores the confidence in the results but also serves as a valuable tool for enhancing training sets. As Liu et al. [109] have shown, this uncertainty information can be leveraged to enrich training data, potentially using generative AI models, thereby improving overall model performance.
@@ V-B1 Acoustic and Industrial Ultrasonic Imaging for Defect Detection and Characterisation
As industry moves towards digitization, the resulting large datasets make ML-based inspection techniques increasingly relevant. Although useful and cost-effective for automating routine NDT processes, the use of ML is nascent in key industrial areas. Notably, in industrial ultrasonic inspection, most ML methods prioritize data-driven over physics-based models [69, 70]. In the NDT supervised method, Ryu et al. [5] proposed a regression-based multilayer perceptron, which was formulated to estimate the plastic properties (e.g., yield strength) of aluminum alloys, utilizing data from ultrasonic tests. Wang et al. utilized CNN for an ultrasonic framework to accurately pinpoint corrosion damage dimensions and locations, despite some forecasting inaccuracies [135]. They also suggested reconstructing sparse signals from limited transducers through compressed sensing and CNN-based deep learning [4]. In addition, a semi-supervised anomaly detection model was introduced to establish a baseline anomaly detection score for the ultrasonic testing dataset [136]. Authors made adjustments to these networks to enhance their efficiency in addressing specific challenges. Shukla et al. [137] present a physics-informed neural network (PINN) to detect surface-breaking cracks in metal plates based on ultrasonic data. Compared with supervised methods, there is very limited literature available on semi-supervised and unsupervised methods for NDT. Luiken et al. did use a self-supervised denoiser to develop a method for flexible ultrasonic data acquisition, enhancing speed while preserving image quality [138].
@@
While numerous studies in ML-based NDT have focused on defect detection, Rao et al. [6] develop a novel approach. Their study introduces a quantitative ultrasonic imaging technique using a fully-supervised network to identify high-contrast defects in multi-layered bonded structures. This technique employs the fully-connected network (FCN) to convert ultrasonic data into longitudinal wave velocity models, which are then used to reconstruct defects in new data sets. In their experimental setup, as shown in the left panel of Fig. 7(a), a CTS-PA322T portable ultrasonic phased array flaw detector was utilized alongside a linear phased array. The phased array had a central frequency of 5 MHz and a pitch of 0.6 mm. The specimen design is detailed in the central panel of Fig. 7(a). The right panel of the same figure displays the reconstruction of this specimen using experimental data. The results demonstrate a high degree of accuracy in identifying the locations and shapes of two notches and disbonds. The reconstructed quantitative information closely matches the known specimen configuration, with artifacts being nearly undetectable.
@@ V-B2 Acoustic temperature imaging in high explosives
Greenhall et al. [139] develop a fully-supervised, simulation/measurement-driven temperature imaging technique that measures temperature distribution directly from the acoustic measurements using a CNN and a hybrid experiment/simulation training process illustrated in Fig. 7(b). The temperature imaging technique was demonstrated on cylindrical containers filled with HE (pentolite 50/50) as they were heated externally from ambient until ignition. During heating, 16 acoustic transducers around the container circumference transmitted/received (Tx/Rx) acoustic bursts centered at 350 kHz and two thermocouples measured temperatures at the wall and center to collect temperatures for training data (Fig. 7(b, Left Panel)).
To acquire training data temperature profiles within the container without excessive thermocouples disrupting the heating, the HE melting process was simulated numerically using an axisymmetric finite element model (COMSOL) and a radial temperature profile $T(r)$ was measured (Fig. 7(b, Central Panel)). To account for modeling errors, the simulated temperatures were corrected to agree with experimental thermocouple measurements. Training/testing samples comprised of data: acoustic waveforms $X$ measured from a single transmitter, by $N_{Rx}$ receivers and a label: a radial temperature profile $T$. The samples $(X,T)$ were then used to train/test a CNN (see Greenhall et al. [139] for details).
@@
The temperature imaging technique based on a CNN and hybrid data provides considerable benefits over traditional temperature measurement techniques. CNN temperature imaging is noninvasive and does not require penetrating the container, unlike thermocouples and thermistors;[140] it provides an internal temperature profile, unlike infrared thermography, which is limited to surface temperatures;[141] and it works for multi-phase, attenuating materials, unlike existing physics-based acoustic temperature imaging techniques, which require mounting transducers on the container interior and/or are limited to single-phase materials with low acoustic attenuation[142, 143, 144, 145, 146, 147].
@@ V-C Medical Ultrasound Computed Tomography
USCT offers several advantages over traditional breast imaging methods like mammography, such as higher sensitivity to invasive cancer, cost-effectiveness, and being non-radiative and compression-free [164, 165]. Recently approved by the U.S. FDA for breast cancer screening in women with dense breasts and as a general diagnostic tool, USCT still faces significant algorithmic challenges. Key among these is improving image quality and reducing the lengthy image reconstruction times, which currently require high-performance, possibly GPU-accelerated computers [166, 167]. This computational demand not only limits USCT’s clinical deployment but also raises costs, hindering its adoption in resource-limited settings. To overcome these challenges, innovative ML methods are being explored. Fan et al. [168] propose a primal-dual gradient descent method for efficient data mapping, Stanziola et al. [8] employ loop unrolling for wave equation modeling, and Liu and Almekkawy [7] utilize physics-informed neural networks to determine sound speed.
@@ V-C1 Ultrasound Tomography for Breast Cancer Detection using ring-array systems
The cited works [66, 65, 67, 68] explore a fully supervised simulation/measurement-driven approach, where a post-processing network learns to map preliminary velocity model estimates to high-resolution versions. These studies are distinguished by their use of anatomically and physiologically realistic numerical breast phantoms to train and assess the methods [66] (refer to Appendix E for more details). Two methods [66, 65] use an early-stopped USCT-derived velocity model as input for a U-net architecture, producing visually enhanced yet potentially inaccurate models. Conversely, Jeong et al. [67, 68] show that using diverse input maps like travel time estimates and reflectivity maps leads to more accurate and reliable velocity models, although some fine structures may be erroneously introduced, as the authors acknowledge [169]. Poudel et al. [170] and Li et al. [171] demonstrate a fully supervised approach that addresses image reconstruction artifacts and inaccuracies arising from the mismatch between 3D wave propagation in the physical system and the simpler 2D imaging model often employed in breast USCT imaging.
@@
Here we highlight the work of Lozenski et al. [9] on enhancing USCT (as shown in Fig. 8). They address the training and assessment of learned USCT methods using task-based measurements of image quality. In this work, a fully supervised physics-aware approach is proposed where a CNN mapping measurement data to the corresponding velocity is trained to minimize a specifically designed loss function that incorporates information specific to the clinical task for which the image was taken. To incorporate task-specific information, a weighted sum of the mean square error loss (MSE) in the image domain and a novel task-informed objective was proposed based on features extracted by a CNN-based numerical observer that was pre-trained on a tumor detection-localization task using the reference velocity models and tumor segmentation maps. Computer simulation studies using an idealized 2D USCT imaging system demonstrate that the proposed learning method can achieve a similar visual appearance to the velocity model estimated by USCT but also provide more reliable detection and localization of tumors.
@@
Figure 8: (a) Geometry of the idealized 2D USCT imaging system in Lozenski et al. [9] with blue dots indicating emitter locations, red dots receivers, and green circle the field of view. (b) Receiver operating curve (ROC) for a tumor detection and localization task; area under the curve (AUC) values reported in the legend. The true speed of sound map (c) and estimates obtained by physics-based USCT (d), task-agnostic learned USCT (e), and the task-informed learned USCT method (f) in Lozenski et al. [9]. Tumor segmentation maps of the object (g), and segmentation maps estimated from reconstructed images with the three methods (h,i,j). True detection/localization is indicated in white and false detection in red. Speed of sound images produced by the task-informed method have similar visual or superior appearance (second rows) and provide more reliable lesion localization and detection compared to that achieved with physics-based or task-agnostic learned USCT. In particular, The ROC curve plot shows that the task-informed method can achieve over 20% higher AUC than physics-based USCT or other task-agnostic learned methods.

### v2

@@ V Computational Wave Imaging Applications
CWI is widely used in various scientific and engineering fields. This section provides examples of employing these methods to solve real-world problems. Appendix D includes source codes, packages, and training datasets, offering practical examples and a foundation for readers to address their own CWI challenges.
@@ V-A1 Leakage Detection for Carbon Sequestration
Geologic carbon sequestration (GCS), crucial in climate change mitigation, involves capturing and storing CO2. Seismic imaging is key for early detection of CO2 leakage, essential for project success. Effective leakage detection hinges on identifying leak presence, location, and magnitude. While traditional seismic methods are commonly used, they often lack early detection precision. Deep learning has brought significant improvements. [148] develops a method for direct CO2 leakage quantification from seismic data. [58] introduces a semi-supervised approach targeting small leaks and imbalanced data scenarios. [74] employs a generative model to enhance rare event detectability. Moving beyond indirect inference from seismic data, [149] uses a generative adversarial network for mapping dynamic saturation changes. [150] create a network for direct CO2 saturation mapping from seismic data. [151, 128] combine seismic, electromagnetic, and gravity data for a comprehensive saturation model with uncertainty estimation. [60] links seismic and electromagnetic data for improved CO2 saturation modeling.
@@
The Kimberlina dataset, generated by the National Risk Assessment Partnership project of the U.S. Department of Energy, is based on the geologic structure of a commercial-scale GCS reservoir at the Kimberlina site in the southern San Joaquin Basin, CA, USA. The simulation includes four modules: a CO2 storage reservoir model (Fig. 11(a)), a wellbore leakage model (Fig. 11(b)), multi-phase flow and reactive transport models of CO2/brine migration in aquifers (Fig. 11(c)), and a geophysical model.
The P-wave velocity maps used in this study, which form part of the geophysical model, are constructed from the stratigraphic and lithologic information at the Kimberlina GCS site (Fig. 11(b)) [146, 147]. Fig. 11(d) shows a set of simulated velocity maps across 20 time steps spanning 200 years. To assess different inversion strategies, Fig. 11(e) compares the reconstructed velocity fields using three methods: physics-based seismic full waveform inversion (FWI) [45], a fully supervised learning approach [51], and a semi-supervised learning method incorporating adaptive data augmentation [58].
The results highlight the benefits of the adaptive augmentation strategy, which can be applied iteratively to refine CWI reconstructions. In each iteration, the current model generates approximate velocity maps from the measured waveform data, which are then used to synthesize new labeled waveform-velocity pairs via forward modeling. These new samples are added to the training set, and the model is retrained. This cycle is repeated, with each round progressively tailoring the training data to the target recorded waveform, thereby improving inversion accuracy. As shown in Fig. 11(e), this process helps better localize and delineate the CO2 leakage plume, particularly in its early stages. Additional implementation details of the adaptive augmentation procedure are provided in Appendix C. In summary, the physics-based FWI lacks the resolution to capture small-scale leakage features; the fully supervised method suffers from shape distortion; whereas the semi-supervised approach delivers more accurate and physically consistent reconstructions.
@@
Figure 12: Illustration of seismic full-waveform inversion with uncertainty quantification applying to noisy data. Ground Truth (Row 1); Results obtained using InversionNet [51] on clean data (Row 2); Results obtained using InversionNet [51] on noisy data with SNR=10 (Row 3); Results obtained using InvNet_UQ [127] on noisy data with SNR=10 (Row 4); Uncertainty map obtained using [127] (Row 5).
@@ V-A2 Seismic Full-Waveform Inversion with Uncertainty Quantification
Robustness to noise is essential for CWI and broader inversion methods. It is important to distinguish “adversarial noise”—deliberate perturbations meant to deceive ML models—from “measurement noise,” which naturally occurs during data collection. Since the latter is far more relevant to CWI, we consider adversarial scenarios to be beyond the scope of this review. In our second seismic FWI example, we specifically address measurement noise. To improve ML models’ resistance to this, researchers have developed strategies such as UQ [127], advanced physics-guided regularization [11], and training with noisy data. We focus on the UQ method InvNet_UQ by [127].
The effectiveness of InvNet_UQ is demonstrated in Fig. 12, where it is applied to data contaminated with noise at a Signal-to-Noise Ratio (SNR) of 10. This is compared to results from InversionNet [51], revealing a significant performance decline for InversionNet under noisy conditions. In contrast, InvNet_UQ maintains higher accuracy in its predictions, showcasing its superior robustness to noise.
In addition to improved prediction quality, InvNet_UQ also generates uncertainty maps, as shown in the bottom row of Fig. 12. These maps quantify the spatial confidence of the inversion output and serve multiple purposes. They can help practitioners identify regions of high uncertainty, informing risk-aware decision-making in geophysical and subsurface applications. Furthermore, these uncertainty estimates can guide the adaptive selection of training data or even act as a supervisory signal to train auxiliary models. Beyond interpretability, the ability to model predictive uncertainty also contributes directly to more robust inversion results. By estimating a conditional distribution over possible outputs, rather than a single point estimate, InvNet_UQ mitigates the risk of overfitting to noise or making overconfident errors in ill-posed regions. This allows the model to “hedge” its predictions in ambiguous areas, producing smoother and more stable reconstructions. The uncertainty maps shown in Fig. 12, calculated using Eq. (14), thus provide not only interpretability but also a mechanism to improve the inversion pipeline itself. As [127] have shown, this uncertainty information can be leveraged to enrich training data, potentially using generative AI models, thereby improving overall model performance.
@@ V-B1 Acoustic and Industrial Ultrasonic Imaging for Defect Detection and Characterisation
As industry moves towards digitization, the resulting large datasets make ML-based inspection techniques increasingly relevant. Although useful and cost-effective for automating routine NDT processes, the use of ML is nascent in key industrial areas. Notably, in industrial ultrasonic inspection, most ML methods prioritize data-driven over physics-based models [79, 80]. In the NDT supervised method, [5] proposes a regression-based multilayer perceptron, which was formulated to estimate the plastic properties (e.g., yield strength) of aluminum alloys, utilizing data from ultrasonic tests. [152] utilizes CNN for an ultrasonic framework to accurately pinpoint corrosion damage dimensions and locations, despite some forecasting inaccuracies. They also suggested reconstructing sparse signals from limited transducers through compressed sensing and CNN-based deep learning [4]. In addition, a semi-supervised anomaly detection model was introduced to establish a baseline anomaly detection score for the ultrasonic testing dataset [153]. Authors made adjustments to these networks to enhance their efficiency in addressing specific challenges. [154] presents a physics-informed neural network (PINN) to detect surface-breaking cracks in metal plates based on ultrasonic data. Compared with supervised methods, there is very limited literature available on semi-supervised and unsupervised methods for NDT. [155] uses a self-supervised denoiser to develop a method for flexible ultrasonic data acquisition, enhancing speed while preserving image quality.
@@
While numerous studies in ML-based NDT have focused on defect detection, [6] develops a novel approach. Their study introduces a quantitative ultrasonic imaging technique using a fully-supervised network to identify high-contrast defects in multi-layered bonded structures. This technique employs the fully-connected network (FCN) to convert ultrasonic data into longitudinal wave velocity models, which are then used to reconstruct defects in new data sets. In their experimental setup, as shown in the left panel of Fig. 13(a), a CTS-PA322T portable ultrasonic phased array flaw detector was utilized alongside a linear phased array. The phased array had a central frequency of 5 MHz and a pitch of 0.6 mm. The specimen design is detailed in the central panel of Fig. 13(a). The right panel of the same figure displays the reconstruction of this specimen using experimental data. The results demonstrate a high degree of accuracy in identifying the locations and shapes of two notches and disbonds. The reconstructed quantitative information closely matches the known specimen configuration, with artifacts being nearly undetectable.
@@ V-B2 Acoustic temperature imaging in high explosives
[156] develops a fully-supervised, simulation/measurement-driven temperature imaging technique that measures temperature distribution directly from the acoustic measurements using a CNN and a hybrid experiment/simulation training process illustrated in Fig. 13(b). The temperature imaging technique was demonstrated on cylindrical containers filled with HE (pentolite 50/50) as they were heated externally from ambient until ignition. During heating, 16 acoustic transducers around the container circumference transmitted/received (Tx/Rx) acoustic bursts centered at 350 kHz and two thermocouples measured temperatures at the wall and center to collect temperatures for training data (Fig. 13(b, Left Panel)).
To acquire training data temperature profiles within the container without excessive thermocouples disrupting the heating, the HE melting process was simulated numerically using an axisymmetric finite element model (COMSOL) and a radial temperature profile $T(r)$ was measured (Fig. 13(b, Central Panel)). To account for modeling errors, the simulated temperatures were corrected to agree with experimental thermocouple measurements. Training/testing samples comprised of data: acoustic waveforms $X$ measured from a single transmitter, by $N_{Rx}$ receivers and a label: a radial temperature profile $T$. The samples $(X,T)$ were then used to train/test a CNN (see [156] for details).
@@
The temperature imaging technique based on a CNN and hybrid data provides considerable benefits over traditional temperature measurement techniques. CNN temperature imaging is noninvasive and does not require penetrating the container, unlike thermocouples and thermistors;[157] provides an internal temperature profile, unlike infrared thermography, which is limited to surface temperatures;[158] works for multi-phase, attenuating materials, unlike existing physics-based acoustic temperature imaging techniques, which require mounting transducers on the container interior and/or are limited to single-phase materials with low acoustic attenuation[159, 160, 161, 162, 163, 164].
@@ V-C Medical Ultrasound Computed Tomography
USCT offers several advantages over traditional breast imaging methods like mammography, such as higher sensitivity to invasive cancer, cost-effectiveness, and being non-radiative and compression-free [181, 182]. Recently approved by the U.S. FDA for breast cancer screening in women with dense breasts and as a general diagnostic tool, USCT still faces significant algorithmic challenges. Key among these is improving image quality and reducing the lengthy image reconstruction times, which currently require high-performance, possibly GPU-accelerated computers [183, 184]. This computational demand not only limits USCT’s clinical deployment but also raises costs, hindering its adoption in resource-limited settings. To overcome these challenges, innovative ML methods are being explored. [185] proposes a primal-dual gradient descent method for efficient data mapping, [8] employs loop unrolling for wave equation modeling, and [7] utilizes physics-informed neural networks to determine sound speed.
@@ V-C1 Ultrasound Tomography for Breast Cancer Detection using ring-array systems
The cited works [76, 75, 77, 78] explore a fully supervised simulation/measurement-driven approach, where a post-processing network learns to map preliminary velocity model estimates to high-resolution versions. These studies are distinguished by their use of anatomically and physiologically realistic numerical breast phantoms to train and assess the methods [76] (refer to Appendix D for more details). Two methods [76, 75] use an early-stopped USCT-derived velocity model as input for a U-net architecture, producing visually enhanced yet potentially inaccurate models. Conversely, [77, 78] show that using diverse input maps like travel time estimates and reflectivity maps leads to more accurate and reliable velocity models, although some fine structures may be erroneously introduced, as the authors acknowledge [186]. [187] and [188] demonstrate a fully supervised approach that addresses image reconstruction artifacts and inaccuracies arising from the mismatch between 3D wave propagation in the physical system and the simpler 2D imaging model often employed in breast USCT imaging.
@@
Here we highlight the work of [9] on enhancing USCT (as shown in Fig. 14). They address the training and assessment of learned USCT methods using task-based measurements of image quality. In this work, a fully supervised physics-aware approach is proposed where a CNN mapping measurement data to the corresponding velocity is trained to minimize a specifically designed loss function that incorporates information specific to the clinical task for which the image was taken. To incorporate task-specific information, a weighted sum of the mean square error loss (MSE) in the image domain and a novel task-informed objective was proposed based on features extracted by a CNN-based numerical observer that was pre-trained on a tumor detection-localization task using the reference velocity models and tumor segmentation maps. Computer simulation studies using an idealized 2D USCT imaging system demonstrate that the proposed learning method can achieve a similar visual appearance to the velocity model estimated by USCT but also provide more reliable detection and localization of tumors.
@@
Figure 14: (a) Geometry of the idealized 2D USCT imaging system in [9] with blue dots indicating emitter locations, red dots receivers, and green circle the field of view. (b) Receiver operating curve (ROC) for a tumor detection and localization task; area under the curve (AUC) values reported in the legend. The true speed of sound map (c) and estimates obtained by physics-based USCT (d), task-agnostic learned USCT (e), and the task-informed learned USCT method (f) in [9]. Tumor segmentation maps of the object (g), and segmentation maps estimated from reconstructed images with the three methods (h,i,j). True detection/localization is indicated in white and false detection in red. Speed of sound images produced by the task-informed method have similar visual or superior appearance (second rows) and provide more reliable lesion localization and detection compared to that achieved with physics-based or task-agnostic learned USCT. In particular, The ROC curve plot shows that the task-informed method can achieve over 20% higher AUC than physics-based USCT or other task-agnostic learned methods.